OpenStudy (anonymous):

How do you solve the following equation by factoring? c^2=5c

OpenStudy (anonymous):

c=0 c=5

OpenStudy (anonymous):

Thanks but can you show how you got that answer?

OpenStudy (anonymous):

c^2 =5c (c)x(c) =5c [(c)x(c)]/(c) = 5c/c c = 5

OpenStudy (anonymous):

Rewrite the equation: c^2 - 5c = 0. Factor out a c: C ( C - 5 ) = 0. What values of C make this true?